$SPAD/src/input subset.input

Timothy Daly December 3, 2016

Abstract

Contents

2

— * —

)set break resume
)sys rm -f subset.output
)spool subset.output
)set message test on
)set message auto off
)clear all
--S 1 of 6
SI ==> Set Integer
--R
--R                                                                   Type: Void
--E 1
--S 2 of 6
SSI ==> Set SI
--R
--R                                                                   Type: Void
--E 2
--S 3 of 6
si:SI := set [2]
--R
--R
--R   (3)  {2}
--R                                                           Type: Set(Integer)
--E 3
--S 4 of 6
subset?(si,si)
--R
--R
--R   (4)  true
--R                                                                Type: Boolean
--E 4
--S 5 of 6
ssi:SSI := set [si]
--R
--R
--R   (5)  {{2}}
--R                                                      Type: Set(Set(Integer))
--E 5
--S 6 of 6
subset?(ssi,ssi)
--R
--R

--R   (6)  true
--R                                                                Type: Boolean
--E 6
)spool
)lisp (bye)

———-

References^{[1] nothing}

5